OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. -15, -6, 3, 12, ... an = -15 + 9(n + 1) an = -15 x 9(n - 1) an = -15 + 9(n + 2) an = -15 + 9(n - 1)

OpenStudy (anonymous):

if n = 1, what do each of those expressions give you?

OpenStudy (anonymous):

actually, start with n = 0 it'll give you the answer right off the bat

OpenStudy (anonymous):

I got A?

OpenStudy (anonymous):

wait. C? @AllTehMaffs

OpenStudy (anonymous):

nevermind. I got it. it was A. thank you :) but would I go about doing the same thing for this one? Find an equation for the nth term of the arithmetic sequence. a14 = -33, a15 = 9 an = -579 + 42(n + 1) an = -579 + 42(n - 1) an = -579 - 42(n + 1) an = -579 - 42(n - 1)

OpenStudy (anonymous):

yeah, just plug in "14" and "15" into the given functions to see which come out as -33 for n=14 and 9 for n=15 ^_^